# Joint Transmit and Pinching Beamforming for Pinching Antenna Systems (PASS): Optimization-Based or Learning-Based?

### v2

@@ VI Conclusion
A novel PASS-enabled multi-user MISO framework has been proposed, which enables pinching beamforming to adjust both large-scale path loss and phases of radiated signals. A joint transmit and pinching beamforming optimization problem was formulated to maximize the system sum rate. To tackle this highly coupled and nonconvex problem, both optimization-based and learning-based methods have been proposed. The optimization-based method, named MM-PDD algorithm, used a convex surrogate function to handle complex exponential components and invoked PDD for problem decoupling, which was guaranteed to obtain stationary solutions. The learning-based method, termed KDL-Transformer, directly reconstructed KKT-conditioned solutions in a low-complexity and data-driven way, where the underlying inter-PA/inter-user and CSI-beamforming dependencies were effectively learned by Transformer. Numerical results have been provided to verify the efficiency of both the proposed PASS framework and algorithms. In future, the design of joint transmit and pinching beamforming under imperfect CSI can be further investigated.
@@
Using the first-order optimality, the equalizer of MMSE receiver $v_{k}^{opt}$ can be obtained by (13). Substituting (13) into (III-A), the MSE $e_{k}^{\mathrm{MSE}}$ is rewritten as
@@
[…]
@@
The complex number $c$ can be written as $c=|c|e^{-i\phi_{c}}$. Using Euler formula $e^{i\theta}=\cos(\theta)+i\sin(\theta)$, we can rewrite function $-\mathrm{Re}\left(ce^{i\left(a\theta\right)}\right)$ into
@@
$\begin{split}\left\|\nabla_{\theta}f\!\left(\theta_{1}\right)\!-\!\nabla_{\theta}f\!\left(\theta_{2}\right)\!\right\|&\leqslant\bigg{(}\max_{\theta\in\left[\theta_{1},\theta_{2}\right]}\left|\nabla_{\theta}^{2}f(\theta)\right|\bigg{)}\left\|\theta_{1}-\theta_{2}\right\|\\ &\leqslant a^{2}\left|c\right|\left\|\theta_{1}-\theta_{2}\right\|,~{}\forall\theta_{1},\theta_{2}\in\mathbb{R}.\end{split}$ (41)

### v3

@@ VI Conclusion
A novel PASS-enabled multi-user MISO framework has been proposed, which enables pinching beamforming to adjust both large-scale path loss and phases of radiated signals. A joint transmit and pinching beamforming optimization problem was formulated to maximize the system sum rate. To tackle this highly coupled and nonconvex problem, both optimization-based and learning-based methods have been proposed. The optimization-based method, named MM-PDD algorithm, used a convex surrogate function to handle complex exponential components and invoked PDD for problem decoupling, which was guaranteed to obtain stationary solutions. The learning-based method, termed KDL-Transformer, directly reconstructed KKT-conditioned solutions in a low-complexity and data-driven way, where the underlying inter-PA/inter-user and CSI-beamforming dependencies were effectively learned by Transformer. Numerical results have been provided to verify the efficiency of both the proposed PASS framework and algorithms. In future, the design of joint transmit and pinching beamforming under practical LoS blockages and waveguide attenuation can be further investigated. Moreover, recent advances in accelerated learning, such as multithreading and federated reinforcement learning [30, 31], can further reduce training overhead and improve real-time adaptability in PASS systems. Prioritized information bottleneck frameworks for distributed online learning [32] also support low-latency and adaptive decision making.
@@
Using the first-order optimality, the equalizer of MMSE receiver $v_{k}^{opt}$ can be obtained by (13). Substituting (13) into (11), the MSE $e_{k}^{\mathrm{MSE}}$ is rewritten as
@@
$\max_{\mathbf{x},\mathbf{D}}~\sum_{k=1}^{K}\left(\log_{2}\left(e_{k}^{\mathrm{MSE}}\right)^{-1}\right),\quad\text{s.t.}~\text{(C1) - (C3)}.$ (38)
@@
The complex number $c$ can be written as $c=|c|e^{-i\phi_{c}}$. Using Euler formula $e^{i\theta}=\cos(\theta)+i\sin(\theta)$, we can rewrite function $\mathrm{Re}\left(ce^{i\left(a\theta\right)}\right)$ into
@@
$\begin{split}\hskip-10.00002pt\left\|\nabla_{\theta}f\!\left(\theta_{1}\right)\!-\!\nabla_{\theta}f\!\left(\theta_{2}\right)\!\right\|&\leqslant\bigg(\max_{\theta\in\left[\theta_{1},\theta_{2}\right]}\left|\nabla_{\theta}^{2}f(\theta)\right|\bigg)\left\|\theta_{1}-\theta_{2}\right\|\\ &\leqslant a^{2}\left|c\right|\left\|\theta_{1}-\theta_{2}\right\|,~\forall\theta_{1},\theta_{2}\in\mathbb{R}.\end{split}$ (41)
@@
-C Proof of Lemma 3
The convergence can be established by combining the descent AL updates and the PDD theory  [19]. (i) AL objective descent. Each inner iteration $t$ builds a tight MM surrogate $f_{\mathrm{MM}}(\bm{\nu};\bm{\nu}^{(t)})\!\geq\!f_{\mathrm{AL}}(\bm{\nu})$ with equality at $\bm{\nu}\!=\!\bm{\nu}^{(t)}$. The BCD update yields $f_{\mathrm{MM}}(\bm{\nu}^{(t+1)};\bm{\nu}^{(t)})\!\leq\!f_{\mathrm{MM}}(\bm{\nu}^{(t)};\bm{\nu}^{(t)})$, hence $f_{\mathrm{AL}}\big(\bm{\nu}^{(t+1)}\big)\!\leqslant\!f_{\mathrm{MM}}\big(\bm{\nu}^{(t+1)};\bm{\nu}^{(t)}\big)\!\leqslant\!f_{\mathrm{MM}}\big(\bm{\nu}^{(t)};\bm{\nu}^{(t)}\big)\!=\!f_{\mathrm{AL}}\big(\bm{\nu}^{(t)}\big)$ and thus $\{f_{\mathrm{AL}}\big(\bm{\nu}^{(t)}\big)\}$ yields a non-increasing sequence. (ii) Robinson condition and equality residuals. The equality constraints (1)-(2) have block-diagonal Jacobians with nonzero diagonal blocks (full row rank), and inequality constraints (C1)-(C3) admit strictly feasible points. Thus, the sufficient condition in [19, Sec. V-A] for Robinson’s condition holds for (P2). Let $\bm{\xi}^{(i)}\!\triangleq\!\bm{\lambda}^{(i)}\!+\!\tfrac{1}{\rho^{(i)}}\mathbf{B}^{(i)}$ at the outer interation $i$. If $\|\mathbf{B}^{(i)}\|_{\infty}\!\leqslant\!\varepsilon^{(i)}$, the dual update $\bm{\lambda}^{(i+1)}\!=\!\bm{\xi}^{(i)}$ gives $\mathbf{B}^{(i)}\!=\!\rho^{(i)}(\bm{\lambda}^{(i+1)}\!-\!\bm{\lambda}^{(i)})$, hence both $\|\mathbf{B}^{(i)}\|_{\infty}$ and $\|\bm{\lambda}^{(i+1)}\!-\!\bm{\lambda}^{(i)}\|_{\infty}$ vanish as $\varepsilon^{(i)}\!\to\!0$. Otherwise $\rho^{(i+1)}\!=\!\varsigma\rho^{(i)}$ with $\varsigma\!\in\!(0,1)$. If the residuals did not decrease, then $\|\bm{\xi}^{(i)}\|\!=\!\|\bm{\lambda}^{(i)}+\tfrac{1}{\rho^{(i)}}\mathbf{B}^{(i)}\|$ would blow up as $\rho^{(i)}\!\to\!0$, contradicting the boundedness of $\{\bm{\xi}^{(i)}\}$ ensured under Robinson’s conditions [19]. Thus, $\|\mathbf{B}^{(i)}\|_{\infty}\!\to\!0$ vanish asymptotically.
Combining (i) and (ii), the inner-loop MM updates ensure that each AL subproblem is solved to an approximate stationary point with a diminishing approximation error, while the outer-loop PDD updates enforce residuals $\|\mathbf{B}^{(i)}\|_{\infty}$ to zero and keep $\{\bm{\lambda}^{(i)}\}$ bounded. As $i\to\infty$, the iterates $(\bm{\nu}^{(i)},\bm{\lambda}^{(i)})$ asymptotically satisfy exact KKT conditions of (P2) with vanishing approximation errors  [19, Theorem 3.1]. Since (P2) is an equivalent reformulation of (P0), the final solution $\bm{\nu}^{\star}$ is a KKT point of (P0), which completes the proof.